OpenStudy (hihi67):

(-x^2)^3/4z ÷ 3x/8z^2 =

OpenStudy (anonymous):

that was wrong i forgot the square

OpenStudy (hihi67):

ooh

OpenStudy (hihi67):

6x/8y^3 ÷ (2x^2)^3/9x^4 =

OpenStudy (anonymous):

start with \[\frac{-x^6\times 8z^2}{4z\times 3x}\] that should work

OpenStudy (anonymous):

cancel a 4, and x and a z to get \[-\frac{2x^5z}{3}\]

OpenStudy (anonymous):

start with \[\frac{6x\times 9x^4}{8y^3\times 8x^6}\]

OpenStudy (hihi67):

sorry. i wrote the problem wrong. it is 9y^4

OpenStudy (hihi67):

and how did you get 8x^6 out of (2x^2)^3

OpenStudy (anonymous):

then start with \[\frac{6x\times 9y^4}{8y^3\times 8x^6}\]

OpenStudy (anonymous):

\[(2x^2)^3=2^3\times (x^2)^3=8\times x^{2\times 3}=8x^6\]

OpenStudy (hihi67):

(2x^2)^3 = 2x^2 * 2x^2 * 2x^2. and 2 * 2 * 2 = 8. oh i see. someone else told me differently.

OpenStudy (anonymous):

same thing right?

OpenStudy (hihi67):

yup

OpenStudy (anonymous):

still not done \[\frac{6x\times 9y^4}{8y^3\times 8x^6}\] \[\frac{27y}{32x^5}\] by canceling

OpenStudy (hihi67):

okay.

OpenStudy (hihi67):

and (a^2)^2/3b / 4a^5/(b^3)^2=

OpenStudy (hihi67):

(a^2)^2/3b / 4a^5/(b^3)^2=